OpenStudy (anonymous):

Giving medal for the answer, plz help ! not good at this Write a rule for the linear function whose graph has slope m and contains the given point. m = 3; (-2, -2) a. y = 3x + 4 c. y = 3x + 2 b. y = 3x - 4 d. y = 3x - 2

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is \(y = mx + b\) where m = the slope, and b = the y-intercpet.

OpenStudy (nurali):

y-y1=m(x-x1) x1=-2 and y1=-2 y+2=3(x+2) y+2=3x+6 y=3x+6-2 y=3x+4

OpenStudy (mathstudent55):

Start writing your equation in that form using the info you have, m = 3: \(y = 3x + b\) To find b, insert the point you are given, (-2, -2), into x and y, and solve for b.

OpenStudy (anonymous):

thanks for the help!

OpenStudy (nurali):

Anytime.

OpenStudy (mathstudent55):

\( -2 = 3(-2) + b \) \(-2 = -6 + b\) \(4 = b\) So you get: \(y = 3x + 4\)